OpenStudy (anonymous):

IVP: y+x(y+1)dy/dx=0 y(e^2)=1

OpenStudy (anonymous):

I just need hints

OpenStudy (jamesj):

c'mon. The first thing to always check is if it is separable. Is it?

OpenStudy (anonymous):

yes, looks like this i think: xdx=-y/(y+1)dy

OpenStudy (jamesj):

double check that

OpenStudy (anonymous):

my algebra is shot, i'm sorry I can"t figure out what it is :(

OpenStudy (jamesj):

in which case you are deep trouble. Solving differential equations assumes that you are comfortable with basic algebraic manipulation. Sorry to be blunt, but don't give up so easily. Try again.

OpenStudy (anonymous):

It's no problem, I know I'm not where I need to be, I'll give it another go

OpenStudy (anonymous):

How about: (y+1)/-y dy= x dx ??? Am I even close?

OpenStudy (anonymous):

\[\Huge \checkmark \] Don't mean to interrupt, I was just looking in.

OpenStudy (anonymous):

Okay :) does ✓ mean that I finally got it right?

OpenStudy (jamesj):

If y+x(y+1)dy/dx=0 then x(y+1) . dy/dx = -y -(y+1)/y . dy/dx = 1/x -(y+1)/y . dy = 1/x . dx

OpenStudy (anonymous):

The problem will be the integration, not the separation, in my opinion.

OpenStudy (anonymous):

Oh I see

OpenStudy (anonymous):

What confuses me entirely is this part: \[\Large \int -\frac{y+1}{y}dy= \int \frac{1}{x}dx \\ \Large \int \left(-1 - \frac{1}{y} \right)dy= \int \frac{1}{x}dx\] so, considering that I didn't sleep to much but I end up with \[\Large -y(x) - \ln |y|=\ln|x| + C \] Can somebody show me where I went wrong?

OpenStudy (anonymous):

That's what I got....

OpenStudy (jamesj):

yes, looks right. Now apply the initial condition and see what you end up with.

OpenStudy (anonymous):

I see, I am used to get the equation in explicit form first, here that would be rather hard it seems.

OpenStudy (jamesj):

Yes, it won't be possible write down an expression for y(x) explicitly, but only implicitly.

OpenStudy (anonymous):

C=-3 ?

OpenStudy (jamesj):

clearly yes

OpenStudy (anonymous):

Woot, thanks a lot for your help James, you've been so helpful!

OpenStudy (jamesj):

I would write this now as y + ln y = - ln x -3 or ln(y.e^y) = ln(1/(e^3.x)) Hence y.e^y = 1/(e^3.x)

OpenStudy (jamesj):

You should now verify this satisfies the IVP by differentiating (implicitly) and checking also double checking the initial value.